# Sergei Gukov, Ludmil Katzarkov, Josef Svoboda

### v2

@@ 3. $Z_{0}$ and splice diagrams
In this section we will concentrate on the series $Z_{0}(q)$ defined as sum of all $\widehat{Z}_{\sigma}(q)$ over all $\sigma\in spin^{c}(Y)$:
$Z_{0}(q)=\sum_{\sigma\in spin^{c}(Y)}\widehat{Z}_{\sigma}(q).$
Clearly, one has
(3.1) $Z_{0}(q)=2^{-s}q^{\frac{-3s-\text{Tr}(M)}{4}}\mathcal{L}_{-M}(f(\vec{z}\,)),$
where $F(\vec{z}\,)$ is give by the expansion (2.5) without the projection $\pi_{\sigma}.$
@@
We will prove that $Z_{0}(q)$ can be reconstructed666up to an overall power from less data than the full plumbing graph (Theorem 3.5). The relevant notion is the splice diagram [65, 50, 19], which was originally developed for the study of singularities but proves to be very useful in our context as well. For homology spheres, the plumbing graph can be reconstructed from the splice diagram, but the splice diagram can be used for faster computations of the unique $\widehat{Z}_{\sigma}(q)=Z_{0}(q)$ in this case.
3.1. Splice diagrams
@@
We will recall the notions of maximal splice diagram and splice diagram. The maximal splice diagram [54] is a useful way how to repackage the data of the plumbing graph $\Gamma$. It is directly related to the inverse matrix of the plumbing matrix, which occurs in the formula for $\widehat{Z}_{\sigma}(q)$. Even more importantly, it gives a good grasp of which entries of this matrix are important and which can be suppressed in some situations. This leads to the notion of the splice diagram. Splice diagrams are useful for constructing algebraic equations of singularities from topology; see Section 3.4.
@@
The maximal splice diagram of $Y$ is a graph on the same set of vertices as $\Gamma$ with a weight $w_{ve}$ for each pair $(v,e)$ of an edge $e$ with its endpoint vertex $v$. The weight $w_{ve}$ is given by the determinant of $-M^{ve}$, where $M^{ve}$ is the plumbing matrix of the connected component of $\Gamma{\setminus}e$ not containing $v$.
@@
Let $Y$ be a manifold given by the plumbing graph in Fig. 2. It is a homology sphere. Its maximal splice diagram and splice diagram are shown at Fig. 3.
Figure 2. Plumbing graph of a homology sphere $Y$.
Figure 3. Maximal splice diagram and splice diagram of $Y.$
The plumbing graph and the maximal splice diagram are essentially equivalent as follows from the next theorem [19]. For $v$, $w$ vertices of the maximal splice diagram, we define $N_{ab}$ as the product of all weights adjacent to the shortest path between $v$ and $w$, not lying on the path. For a matrix $A$, denote $\operatorname{adj}(A)$ the adjugate matrix of $A$, i.e. $\operatorname{adj}(A)=\det(A)A^{-1}$.
Theorem 3.3 ([54, Thm. 12.2]).
Let $M$ be a plumbing matrix of a negative definite manifold $Y$ and let $N$ be as above. Then we have
$N=\operatorname{adj}(-M).$
Therefore, the maximal splice diagram of $Y$ determines the plumbing graph $\Gamma$ of $Y$.
@@
The theorem is illustrated on Example 3.2 in Fig. 4.
222255552222vw
Figure 4. The entry $\operatorname{adj}(-M)_{vw}=2\cdot 2\cdot 5$ is given by the product of adjacent weights to the path connecting two vertices.
3.2. Universal abelian covers
Recall our assumption that $Y$ is a rational homology sphere, so $H=H_{1}(Y,\mathbb{Z})$ is finite with corresponding finite universal abelian cover $Y^{ab}.$ Universal abelian covers are related to splice diagrams by the following theorem, explaining their topological significance.
@@ 3.2. Universal abelian covers
It is often useful to consider $Y$ as the quotient of $Y^{ab}$ by $H$ as quantities attached to $Y^{ab}$ are often simpler, and then they can be subsequently equivariantly refined using the action of $H$. In singularity theory, $Y^{ab}$ is sometimes a link of a simpler singularity, e.g. it is of Brieskorn diagonal type for quasi-homogeneous singularities. One can think of $Z_{0}(q)$ as being attached to $Y^{ab}$ and $\widehat{Z}_{\sigma}(q)$ being the $H$-equivariant refinement. This is formalized in Theorem 3.5 and Corollary 3.7.
Note that for $Y$ plumbed, the universal abelian cover $Y^{ab}$ is plumbed as well. It does not have to be, however, a rational homology sphere, when $Y$ is, or some intermediate covering may lack this property. Unlike for the universal cover, taking the universal abelian cover is not an idempotent operation, i.e., as it may have nontrivial $H$ again and the construction can be repeated. For example, one has a tower of links of singularities
[…]
It is an interesting question how the $\widehat{Z}_{\sigma}(q)$ invariants behave with respect to taking covers. The only clue we have is the relation for lens spaces and their universal abelian cover, 3-sphere $S^{3}$, see Section 4.2.
3.3. $Z_{0}$ and splice diagrams
We will now prove that the series $Z_{0}(q)$ can be reconstructed from the splice diagram, up to a prefactor, which is proportional to Casson–Walker invariant $\lambda(Y)$ [70]. Let $H=H_{1}(Y,\mathbb{Z})$ and $\lambda(Y)$ be the Casson–Walker invariant.
@@ 3.3. $Z_{0}$ and splice diagrams
Let $Y$ be a negative definite plumbed manifold which is a rational homology sphere. Then the $q$-series
(3.2) $q^{-6\lambda(Y)}Z_{0}(q^{|H|})$
only depends on the splice diagram of $Y$.
@@
We will express the elements of the inverse matrix using Theorem 3.3. Then we will find that those entries which are not expressible using splice diagram contribute only by an overall power of $q.$ Here the integrand of $Z_{0}(q)$ plays an important role. Finally we will relate this overall power to $\lambda(Y)$.
The series $Z_{0}(q^{|H|})$ can be written, up to prefactor $q^{\Lambda}$, as a sum of terms
(3.3) $q^{-(\vec{l},\det(-M)M^{-1}\vec{l})/4}=q^{(\vec{l},\operatorname{adj}(-M)\vec{l})/4}=\prod_{v,w\in\operatorname{Vert}}q^{l_{v}l_{w}\operatorname{adj}(-M)_{vw}/4}$
Now if we look at the expansion of
(3.4) $(z_{v}-z_{v}^{-1})^{2-\deg(v)}$
we see that the vector $\vec{l}$ has zero components for vertices of degree two. The only entries of $\operatorname{adj}(-M)$ which enter the formula are therefore $\operatorname{adj}(-M)_{vw}$ with $v,w$ leaves or nodes. By Theorem 3.3, all these entries are encoded in the splice diagram, except from $\operatorname{adj}(-M)_{vv}$ for a leaf $v$. In the product (3.3), they contribute by
(3.5) $\prod_{v\in\operatorname{Vert}}q^{\operatorname{adj}(-M)_{vv}l^{2}_{v}/4}=q^{\sum_{v\in\text{Leaves}}\operatorname{adj}(-M)_{vv}/4}$
because for leaves, we have $l_{v}^{2}=(\pm 1)^{2}=1$, from the term $(z_{v}-z_{v}^{-1}).$ We see that these entries contribute to each monomial by the same overall power of $q$. Up to this power, $Z_{0}(q^{|H|})$ depends only on the splice diagram.
@@
(3.6) $\square:=|H|(-3s-\text{Tr}(M))/4+\sum_{v\in\text{Leaves}}\operatorname{adj}(-M)_{vv}/4.$
A well-known formula for Casson–Walker invariant of negative definite plumbings $\lambda(Y)$ reads [47, p. 402]:
(3.7) $-\frac{24}{|H|}\lambda(Y)=\text{Tr}(M)+3s+\sum_{v\in\operatorname{Vert}}(2-\delta_{v})M^{-1}_{vv}$
@@
(3.8) $\square=6\lambda(Y)-\sum_{v\in\text{Nodes}}(2-\delta_{v})\operatorname{adj}(-M)_{vv}/4.$
@@
Using the relation of splice diagrams with universal abelian covers in Theorem 3.4, we obtain the following corollary, relating the series $Z_{0}(q)$ of two non-homeomorphic 3-manifolds:
@@
$q^{-6\lambda(Y_{1})}Z_{0}(Y_{1},q^{|H_{1}(Y_{1})|})=q^{-6\lambda(Y_{2})}Z_{0}(Y_{2},q^{|H_{1}(Y_{2})|}).$
@@
Let us take the manifold from Example 3.2 and call it $Y_{1}$. Another manifold $Y_{2}$ with the same splice diagram and $H_{1}=\mathbb{Z}/17\mathbb{Z}$ is given in Fig. 5.
Figure 5. Two plumbed manifolds with the same splice diagram.
@@
$Z_{0}(Y_{1},q)=\frac{1}{2}q^{7/2}$ $(-1+q-2q^{2}+q^{3}+q^{5}+3q^{9}+q^{10}-q^{14}-q^{16}-q^{17}+\dots$
$Z_{0}(Y_{2},q^{17})=\frac{1}{2}q^{-53/2}$ $(-1+q-2q^{2}+q^{3}+q^{5}+3q^{9}+q^{10}-q^{14}-q^{16}-q^{17}+\dots$
We see that they coincide up to an overall power. The Casson-Walker invariants are $\lambda(Y_{1})=-4$ and $\lambda(Y_{2})=-9,$ respectively.
Note that the series $Z_{0}(Y_{2},q^{17})$ decomposes into 17 $q$-series $\widehat{Z}_{\sigma}(Y_{2},q^{17}).$ Since the $Z_{0}$ series for $Y_{1}$ and $Y_{2}$ coincide up to the overall factor and scaling, we also have a decomposition of $Z_{0}(Y_{1},q)$. This decomposition would be far from obvious without the knowledge of $Y_{2}$ or, equivalently, the $\mathbb{Z}/17\mathbb{Z}$ action on $Y_{1}.$
@@
For rational homology spheres that are not Seifert-fibered, there is always a finite number of manifolds with a given splice diagram. Their number, however, grows rapidly [52, prop. 4.3]. When $Y$ is an integral homology sphere, then the splice diagram determines the maximal splice diagram and hence the plumbing graph [19].
3.4. Splice and splice-quotient singularities
In singularity theory, splice diagrams are used to define splice-quotient singularities, [54]. This important class includes, e.g., weighted homogeneous, and rational singularities. Starting with a plumbed rational homology sphere $Y$ and its splice diagram, one first constructs the equations for the splice-type singularity [^8] More precisely, an equisingular family of singularities. $X^{ab}$. These are equations in variables corresponding to the leaves of the splice diagram. They define an isolated complete intersection surface singularity. The splice-quotient $X$ is the quotient of $X^{ab}$ by a natural action of $H_{1}(Y).$ Note that the construction relies on certain technical conditions on the plumbing graph.
The importance of these singularities stems from their topological nature. They provide a good testing ground for comparing topological and analytical invariants. On the other hand, one should keep in mind that they are very special in some sense, even among the singularities with given link $Y$. A nice illustration of the above philosophy is provided by a theorem of Némethi:
@@ 3.4. Splice and splice-quotient singularities
As we explained in Section 2, $\widehat{Z}_{\sigma}(q)$ can be thought of as $H$-equivariant decomposition of $Z_{0}$. Hence we expect that they should be related to the the splice-quotient $X$, whose link is the manifold $Y.$ The series $Z_{0}(q)$ should be related to the splice singularity, the universal abelian cover $X^{ab}$.
@@
While we cannot provide the answer at the moment, we observe in Theorem 3.5 and Theorem 4.2, that the structural properties of $\widehat{Z}_{\sigma}(q)$ are strikingly similar to the properties of invariants, such as the topological Poincaré series or Seiberg–Witten invariants, which have analytic analogs [47].
4. Seifert manifolds

### v3

@@ 3 $\boldsymbol{Z_{0}}$ and splice diagrams
In this section, we will concentrate on the series $Z_{0}(q)$ defined as the sum of $\widehat{Z}_{\sigma}(q)$ over all $\sigma\in{\rm spin}^{c}(Y)$, $Z_{0}(q)=\sum_{\sigma\in{\rm spin}^{c}(Y)}\widehat{Z}_{\sigma}(q)$. Clearly, one has
$Z_{0}(q)=2^{-s}q^{\frac{-3s-\text{Tr}(M)}{4}}\mathcal{L}_{-M}(F(\vec{z})),$
where $F(\vec{z})$ is given by the expansion (2.3) without the projection $\pi_{\sigma}$.
@@
We will prove that $Z_{0}(q)$ can be reconstructed [^8] Up to an overall power. from less data than the full plumbing graph (see Theorem 3.5). The relevant notion is the splice diagram [19, 50, 64], which was originally developed for the study of singularities but proves to be very useful in our context as well. For homology spheres, the plumbing graph can be reconstructed from the splice diagram, but the splice diagram can be used for faster computations of the unique $\widehat{Z}_{\sigma}(q)=Z_{0}(q)$ in this case.
3.1 Splice diagrams
@@
We will recall the notions of maximal splice diagram and splice diagram. The maximal splice diagram [54] is a useful way to repackage the data of the plumbing graph $\Gamma$. It is directly related to the inverse matrix of the plumbing matrix, which occurs in the formula for $\widehat{Z}_{\sigma}(q)$. Even more importantly, it gives a good grasp of which entries of this matrix are important and which can be suppressed in some situations. This leads to the notion of the splice diagram. Splice diagrams are useful for constructing algebraic equations of singularities from topology; see Section 3.4.
@@
The maximal splice diagram of $Y$ is a graph on the same set of vertices as $\Gamma$ with a weight $w_{ve}$ for each pair $(v,e)$ of an edge $e$ with its endpoint vertex $v$. The weight $w_{ve}$ is given by the determinant of $M^{ve}$, where $M^{ve}$ is the plumbing matrix of the connected component of $\Gamma{\setminus}e$ not containing $v$.
@@
Let $Y$ be a manifold given by the plumbing graph in Figure 2. It is a homology sphere. Its maximal splice diagram and splice diagram are shown at Figure 3.
Figure 2: Plumbing graph of a homology sphere $Y$.
Figure 3: Maximal splice diagram and splice diagram of $Y$.
The plumbing graph and the maximal splice diagram are essentially equivalent as follows from the next theorem [19]. For $v$, $w$ vertices of the maximal splice diagram, we define $N_{vw}$ as the product of all weights adjacent to the shortest path between $v$ and $w$, not lying on the path. For a matrix $A$, denote $\operatorname{adj}(A)$ the adjugate matrix of $A$, i.e., $\operatorname{adj}(A)=\det(A)A^{-1}$.
Theorem 3.3 ([54, Theorem 12.2]).
Let $M$ be a plumbing matrix of a negative definite manifold $Y$ and let $N$ be as above. Then we have $N=\operatorname{adj}(-M)$. Therefore, the maximal splice diagram of $Y$ determines the plumbing graph $\Gamma$ of $Y$.
@@
The theorem is illustrated on Theorem 3.2 in Figure 4.
225522vw
Figure 4: The entry $\operatorname{adj}(-M)_{vw}=2\cdot 2\cdot 5$ is given by the product of adjacent weights to the path connecting two vertices.
3.2 Universal abelian covers
Recall our assumption that $Y$ is a rational homology sphere, so $H=H_{1}(Y,\mathbb{Z})$ is finite with corresponding finite universal abelian cover $Y^{\rm ab}$. Universal abelian covers are related to splice diagrams by the following theorem, explaining their topological significance.
@@ 3.2 Universal abelian covers
It is often useful to consider $Y$ as the quotient of $Y^{\rm ab}$ by $H$ as quantities attached to $Y^{\rm ab}$ are often simpler, and then they can be subsequently equivariantly refined using the action of $H$. In singularity theory, $Y^{\rm ab}$ is sometimes a link of a simpler singularity, e.g., it is of Brieskorn diagonal type for quasi-homogeneous singularities. One can think of $Z_{0}(q)$ as being attached to $Y^{\rm ab}$ and $\widehat{Z}_{\sigma}(q)$ being the $H$-equivariant refinement. This is formalized in Theorem 3.5 and Corollary 3.7.
Note that for $Y$ plumbed, the universal abelian cover $Y^{\rm ab}$ is plumbed as well. It does not have to be, however, a rational homology sphere, when $Y$ is, or some intermediate covering may lack this property. Unlike for the universal cover, taking the universal abelian cover is not an idempotent operation, i.e., as it may have nontrivial $H$ again and the construction can be repeated. For example, one has a tower of links of singularities
$S^{3}$$A_{1}$$D_{4}$$E_{6}$$E_{7}.$$\scriptstyle{2:1}$$\scriptstyle{4:1}$$\scriptstyle{3:1}$$\scriptstyle{2:1}$
It is an interesting question how the $\widehat{Z}_{\sigma}(q)$ invariants behave with respect to taking covers. The only clue we have is the relation for lens spaces and their universal abelian cover, the 3-sphere $S^{3}$, see Section 4.2.
3.3 $\boldsymbol{Z_{0}}$ and splice diagrams
We will now prove that the series $Z_{0}(q)$ can be reconstructed from the splice diagram, up to a prefactor, which is proportional to the Casson–Walker invariant $\lambda(Y)$ [69]. Let $H=H_{1}(Y,\mathbb{Z})$ and $\lambda(Y)$ be the Casson–Walker invariant.
@@ 3.3 $\boldsymbol{Z_{0}}$ and splice diagrams
Let $Y$ be a negative definite plumbed manifold which is a rational homology sphere. Then the $q$-series $q^{-6\lambda(Y)}Z_{0}\bigl{(}q^{|H|}\bigr{)}$ only depends on the splice diagram of $Y$.
@@
We will express the elements of the inverse matrix $M^{-1}$ using Theorem 3.3. Then we will find that those entries which are not expressible using splice diagram contribute only by an overall power of $q$. Here the integrand of $Z_{0}(q)$ plays an important role. Finally, we will relate this overall power to $\lambda(Y)$.
The series $Z_{0}\bigl{(}q^{|H|}\bigr{)}$ can be written, up to prefactor $q^{\Lambda}$, as a sum of terms
$q^{-(\vec{l},\det(-M)M^{-1}\vec{l})/4}=q^{(\vec{l},\operatorname{adj}(-M)\vec{l})/4}=\prod_{v,w\in\operatorname{Vert}}q^{l_{v}l_{w}\operatorname{adj}(-M)_{vw}/4}$ (3.1)
for $\vec{l}\in 2\mathbb{Z}^{s}+\vec{\delta}$. Now if we look at the expansion of
$\bigl{(}z_{v}-z_{v}^{-1}\bigr{)}^{2-\deg(v)}$ (3.2)
we see that a vector $\vec{l}$ contributes to $Z_{0}\bigl{(}q^{|H|}\bigr{)}$ only if it has zero components for vertices of degree two. The only entries of $\operatorname{adj}(-M)$ which enter the formula are therefore $\operatorname{adj}(-M)_{vw}$ with $v$, $w$ leaves or nodes. By Theorem 3.3, all these entries are encoded in the splice diagram, except from $\operatorname{adj}(-M)_{vv}$ for a leaf $v$. In the product (3.1), they contribute by
$\prod_{v\in\text{Leaves}}q^{\operatorname{adj}(-M)_{vv}l^{2}_{v}/4}=q^{\sum_{v\in\text{Leaves}}\operatorname{adj}(-M)_{vv}/4}$ (3.3)
because for leaves, we have $l_{v}^{2}=(\pm 1)^{2}=1$, from the term $\bigl{(}z_{v}-z_{v}^{-1}\bigr{)}$. We see that these entries contribute to each monomial by the same overall power of $q$. Up to this power, $Z_{0}\bigl{(}q^{|H|}\bigr{)}$ depends only on the splice diagram.
@@
$\square:=|H|(-3s-\text{Tr}(M))/4+\sum_{v\in\text{Leaves}}\operatorname{adj}(-M)_{vv}/4.$
A well-known formula for Casson–Walker invariant of negative definite plumbings $\lambda(Y)$ reads [48, p. 296]
$\frac{24}{|H|}\lambda(Y)=\text{Tr}(M)+3s+\sum_{v\in\operatorname{Vert}}(2-\delta_{v})M^{-1}_{vv},$ (3.4)
which can be rewritten using $|H|=\det(-M)$ as
$6\lambda(Y)=|H|(-3s-\text{Tr}(M))/4+\sum_{v\in\operatorname{Vert}}(2-\delta_{v})\operatorname{adj}(-M)_{vv}/4,$ (3.5)
@@
$\square=6\lambda(Y)-\sum_{v\in\text{Nodes}}(2-\delta_{v})\operatorname{adj}(-M)_{vv}/4.$
@@
Using the relation of splice diagrams with universal abelian covers in Theorem 3.4, we obtain the following corollary, relating the series $Z_{0}(q)$ of two non-homeomorphic 3-manifolds.
@@
$q^{-6\lambda(Y_{1})}Z_{0}\bigl{(}Y_{1},q^{|H_{1}(Y_{1})|}\bigr{)}=q^{-6\lambda(Y_{2})}Z_{0}\bigl{(}Y_{2},q^{|H_{1}(Y_{2})|}\bigr{)}.$
@@
Let us take the manifold from Theorem 3.2 and call it $Y_{1}$. Another manifold $Y_{2}$ with the same splice diagram and $H_{1}=\mathbb{Z}/17\mathbb{Z}$ is given in Figure 5.
Figure 5: Two plumbed manifolds with the same splice diagram.
@@
$Z_{0}(Y_{1},q)=\frac{1}{2}q^{7/2}\bigl{(}-1+q-2q^{2}+q^{3}+q^{5}+3q^{9}+q^{10}-q^{14}-q^{16}-q^{17}+\cdots\bigr{)},$
$Z_{0}\bigl{(}Y_{2},q^{17}\bigr{)}=\frac{1}{2}q^{-53/2}\bigl{(}-1+q-2q^{2}+q^{3}+q^{5}+3q^{9}+q^{10}-q^{14}-q^{16}-q^{17}+\cdots\bigr{)}.$
We see that they coincide up to an overall power. The Casson–Walker invariants are $\lambda(Y_{1})=-4$ and $\lambda(Y_{2})=-9,$ respectively.
Note that the series $Z_{0}\bigl{(}Y_{2},q^{17}\bigr{)}$ decomposes into 17 $q$-series $\widehat{Z}_{\sigma}\bigl{(}Y_{2},q^{17}\bigr{)}$. Since the $Z_{0}$ series for $Y_{1}$ and $Y_{2}$ coincide up to the overall factor and scaling, we also have a decomposition of $Z_{0}(Y_{1},q)$. This decomposition would be far from obvious without the knowledge of $Y_{2}$ or, equivalently, the $\mathbb{Z}/17\mathbb{Z}$ action on $Y_{1}$.
@@
For rational homology spheres that are not Seifert-fibered, there is always a finite number of manifolds with a given splice diagram. Their number, however, grows rapidly [52, Proposition 4.3]. When $Y$ is an integral homology sphere, then the splice diagram determines the maximal splice diagram and hence the plumbing graph [19].
3.4 Splice and splice-quotient singularities
In singularity theory, splice diagrams are used to define splice-quotient singularities [54]. This important class includes, e.g., weighted homogeneous, and rational singularities. Starting with a plumbed rational homology sphere $Y$ and its splice diagram, one first constructs the equations for the splice-type singularity [^10] More precisely, an equisingular family of singularities. $X^{\rm ab}$. These are equations in variables corresponding to the leaves of the splice diagram. They define an isolated complete intersection surface singularity. The splice-quotient $X$ is the quotient of $X^{\rm ab}$ by a natural action of $H_{1}(Y)$. Note that the construction relies on certain technical conditions on the plumbing graph.
The importance of these singularities stems from their topological nature. They provide a good testing ground for comparing topological and analytical invariants. On the other hand, one should keep in mind that they are very special in some sense, even among the singularities with given link $Y$. A nice illustration of the above philosophy is provided by a theorem of Némethi.
@@ 3.4 Splice and splice-quotient singularities
As we explained in Section 2, $\widehat{Z}_{\sigma}(q)$ can be thought of as $H$-equivariant decomposition of $Z_{0}$. Hence we expect that they should be related to the splice-quotient $X$, whose link is the manifold $Y$. The series $Z_{0}(q)$ should be related to the splice singularity, the universal abelian cover $X^{\rm ab}$.
@@
While we cannot provide the answer at the moment, we observe in Theorems 3.5 and 4.2, that the structural properties of $\widehat{Z}_{\sigma}(q)$ are strikingly similar to the properties of invariants, such as the topological Poincaré series or Seiberg–Witten invariants, which have analytic analogs [47].
4 Seifert manifolds